### v1

@@ 1.4 Organization of the paper
The remainder of the paper is organized as follows. In Section 2, we present several important preliminaries, including the formal definition of Cannings trees, Aldous’ theorem on the convergence of contour functions, and the construction of the conditioned Brownian motion $W^{\ell}$. Section 3 introduces the coalescent process and derives some of its basic properties, including the coming-down-from-infinity property. Section 4 is devoted to proving the tightness of the rescaled trees. Finally, in Section 5, we complete the proof of the main theorem by establishing the finite-dimensional convergence and the joint convergence of the contour and height functions. Appendix A proves convergence to Kingman’s coalescent under (H2”), and Appendix B provides a counterexample demonstrating the near-sharpness of (H2’).
As a brief guide to the assumptions, our proof of Theorem 1.1 proceeds in two steps. First, under (H1), (H2’) and (H2”), we prove the convergence stated in Theorem 1.1. Second, we show that (H2) implies both (H2’) and (H2”). Accordingly, most results in Sections 35 are stated under (H2’) and (H2”), rather than (H2).
@@ 2.1 Cannings tree
where by convention $(\mathbb{N}_{+})^{0}:=\{\emptyset\}$. Every rooted ordered tree can be canonically represented as a prefix-stable finite subset of $\mathcal{U}$. See, for example, [le2006random, Section 1.1] for details.
Using this formalism, we now give a precise definition of the Cannings tree associated with an inhomogeneous Cannings model. For a sequence $u=(u_{1},\ldots,u_{m})\in\mathcal{U}$, we write $|u|=m$ for its length. For $u=(u_{1},\ldots,u_{m})$ and $j\in\mathbb{N}_{+}$, we denote by $uj$ the concatenated sequence $(u_{1},\ldots,u_{m},j)$. For $u,v\in\mathcal{U}$, we write $u\prec v$ if $u$ precedes $v$ in the lexicographical order on $\mathcal{U}$.
@@
Intuitively, this construction starts by introducing an artificial ancestor, the root node $\emptyset$, which serves as the common parent of all individuals in generation $1$. Each vertex at height $s$ (i.e., with label length $s$) represents an individual in generation $s$, and the corresponding entry $\nu^{s}_{i}$ in the offspring vector determines the number of children that the $i$-th individual (in lexicographic order) in generation $s$ has. These children are then represented by appending $1,2,\ldots,\nu^{s}_{i}$ to the parent’s representation.
In this way, the tree structure encodes the entire genealogy of the population.
@@ 2.2 Foundations of tree encoding and convergence
This section introduces the framework for encoding trees as height and contour functions, and then provides necessary and sufficient conditions for the convergence of contour functions.
@@ 2.2.1 Geometric trees
In this section, we extend the notion of rooted ordered trees by assigning a positive length to each edge, thereby introducing the concept of a rooted ordered geometric tree. This enriched structure is better suited for studying convergence, where edge lengths need to be rescaled.
@@
We denote the root of $T$ by $\rho$. The function $w$ naturally induces a metric on $V(T)$. For any two vertices $v,v^{\prime}\in V(T)$, the distance is given by
$d(v,v^{\prime}):=\sum_{e\in[\![v,v^{\prime}]\!]}w(e),$
where $v,v^{\prime}]\!]$ denotes the unique path connecting $v$ and $v^{\prime}$ in $T$. In particular,
@@
The height function $H(T)$ of a rooted ordered tree $(T,w)$ records the height of each vertex according to the lexicographic order on $T$. Specifically, if
$V(T)=\{v_{0},v_{1},\dots,v_{k-1}\}\quad\text{with}\quad\rho=v_{0}\prec v_{1}\prec\cdots\prec v_{k-1},$
then the height function is defined by
@@ 2.2.2 Encoding compact real trees by bridge functions
Following Le Gall’s survey [jean2005random], compact real trees are defined as compact metric spaces satisfying the standard tree properties (see [jean2005random, Section 2.1] for details). A canonical way to construct such trees, which simultaneously endows them with a root and an ordered structure, is through encoding functions drawn from the space
@@
Specifically, given $g\in C_{\rm bridge}$, we define a pseudo-metric:
$d_{g}(s,t):=g(s)+g(t)-2\inf_{r\in[s\wedge t,s\vee t]}g(r),\quad\text{for }s,t\geq 0.$
This induces an equivalence relation:
@@
The quotient space $\mathcal{T}_{g}:=[0,1]/\sim$, equipped with the inherited metric $d_{g}$, is a compact real tree by [jean2005random, Theorem 2.2]. Let $p_{g}:[0,1]\to\mathcal{T}_{g}$ be the canonical projection. We regard $\mathcal{T}_{g}$ as rooted at $p_{g}(0)$ and equip it with an order by declaring $s_{1}\prec_{g}s_{2}$ if and only if
@@
Le Gall’s encoding of compact real trees uses continuous functions with compact support [jean2005random]. Our encoding through $C_{\rm bridge}$ functions is essentially equivalent because, in Le Gall’s setting, the function $g(x)$ and its rescaled version $g(cx)$ encode the same real tree. Thus, without loss of generality, we can always assume $g\in C_{\rm bridge}$.
@@
where $x,y]\!]\subset\mathcal{T}_{g}$ denotes the unique geodesic path between $x,y\in\mathcal{T}_{g}$.
This subtree is naturally viewed as a rooted ordered geometric tree: its vertex set consists of the root $\rho$, the sampled points $p_{g}(t_{1}),\dots,p_{g}(t_{k})$, and all branch points along their geodesics; edges are the geodesic segments connecting consecutive vertices, endowed with their metric lengths; and the left–right order is inherited from $\preceq_{g}$.
@@
Suppose $g,g^{\prime}\in C_{\rm bridge}$ differ only by a strictly increasing time change; that is, there exists a strictly increasing homeomorphism $\phi:[0,1]\to[0,1]$ such that
$g(t)=g^{\prime}\big(\phi_{t}\big),\qquad t\in[0,1].$
Then the map
$\Psi:\mathcal{T}_{g}\longrightarrow\mathcal{T}_{g^{\prime}},\qquad\Psi\big(p_{g}(t)\big):=p_{g^{\prime}}\big(\phi_{t}\big),$
defines an isomorphism between the rooted ordered real trees $\mathcal{T}_{g}$ and $\mathcal{T}_{g^{\prime}}$. In particular, we have for any $t_{1},\ldots,t_{k}\in[0,1]$, the subtrees
$\mathcal{T}_{g}(t_{1},\ldots,t_{k})\quad\text{and}\quad\mathcal{T}_{g^{\prime}}(\phi_{t_{1}},\ldots,\phi_{t_{k}})$
are isomorphic.
@@ 2.2.3 Convergence of contour functions
where $b^{i}_{(0)}=b^{i}_{(k)}$ is the root of $T^{i}$, and for $1\leq j\leq k-1$, $b^{i}_{(j)}$ denotes the branch point of $V^{i}_{(j)}$ and $V^{i}_{(j+1)}$ in $T^{i}$.
@@
Let $(\mathcal{T},d)=(\mathcal{T}_{g},d_{g})$ be a random rooted ordered real tree encoded by a random $g\in C_{\rm bridge}$. Consider a sequence $(\mathscr{T}^{n}:n\geq 1)$ of random ordered discrete trees with $a_{n}:=|V(\mathscr{T}^{n})|\to\infty$, and suppose that the maximal edge length of $\mathscr{T}^{n}$ converges to 0 in distribution. Then the following are equivalent:
@@
    For each $1\leq k\leq a_{n}$, let $(V_{n,1},\dots,V_{n,a_{n}})$ be a uniform random ordering of the vertices of $\mathscr{T}^{n}$, and let $\mathscr{T}^{n}_{k}$ be the rooted ordered subtree spanned by the root $\rho$ and $\{V_{n,1},\dots,V_{n,k}\}$. Then it holds that:
    • Tightness: define
      $\Delta(n,k):=\max_{v\in V(\mathscr{T}^{n})}\min_{u\in V(\mathscr{T}^{n}_{k})}d(v,u).$ (5)
      Then, for any $\delta>0$,
      $\lim_{k\to\infty}\limsup_{n\to\infty}\mathbb{P}(\Delta(n,k)>\delta)=0.$
    • Finite-dimensional convergence: for every $k\in\mathbb{N}_{+}$,
      $\big(\mathscr{T}^{n}_{k},\{V_{n,1},\dots,V_{n,k}\}\big)\overset{d}{\longrightarrow}\big(\mathcal{T}_{g,k},\{p_{g}(U_{1}),\dots,p_{g}(U_{k})\}\big)$ (6)
      under the distance (4), where $\{V_{n,1},\dots,V_{n,k}\}$ and $\{p_{g}(U_{1}),\dots,p_{g}(U_{k})\}$ are considered as multisets.
@@
Aldous originally proved this equivalence in [1993Aldous, Theorem 20] for the case where the limiting tree is a compact binary real tree encoded by an excursion. In that setting, his notion of finite-dimensional distributions differs from (6). Nevertheless, the first step of his proof shows that his definition is equivalent to requiring, for every $k\geq 1$,
$\begin{aligned} &\big(C_{U_{n,(1)}}(\mathscr{T}^{n}),\inf_{U_{(1)}<t<U_{(2)}}C_{t}(\mathscr{T}^{n}),C_{U_{n,(2)}}(\mathscr{T}^{n}),\inf_{U_{(2)}<t<U_{(3)}}C_{t}(\mathscr{T}^{n}),\ldots,C_{U_{n,(k)}}(\mathscr{T}^{n})\big)\\ &\overset{d}{\longrightarrow}\big(g(U_{(1)}),\inf_{U_{(1)}<t<U_{(2)}}g(t),g(U_{(2)}),\inf_{U_{(2)}<t<U_{(3)}}g(t),\ldots,g(U_{(k)})\big),\end{aligned}$ (7)
where $(U_{n,1},\ldots,U_{n,k})$ is a random vector sampled uniformly without replacement from $\{0,\ldots,2(a_{n}-1)\}$, with $U_{n,(i)}$ denoting its order statistics, and $U_{(1)},\ldots,U_{(k)}$ are the order statistics of $k$ i.i.d. uniform $(0,1)$ random variables. By [1993Aldous, Lemma 12], (7) is readily seen to be equivalent to (6). This is the only step in his argument that relies on the assumption of a binary limiting tree; the remainder of the proof proceeds without invoking it. Thus, Theorem 2.6 is already implicit in Aldous’ proof.
The formulation in (6) is more suitable for our setting, where the root of the limiting tree is not necessarily binary.
@@ 2.3 Conditioned Brownian motion and its subtree distribution
Recall the notion of a continuous profile introduced in Section 1. In this section, we first construct a canonical version of $W^{\ell}$ based on Aldous’ construction [1998BROWNIAN]. We then study the laws of the uniform $k$-point subtrees induced by $W^{\ell}$ and by the time-changed process $W^{\ell^{\sigma}}_{\alpha^{-1}(t)}$. These results will be used in the proof of finite-dimensional convergence in Section 5.1.
@@ 2.3.1 Canonical version of $W^{\ell}$
Recall that $L_{t}(x)$ denotes the local time of $W$ at level $x$ up to time $t$, and that $\tau_{u}$ represents the inverse local time at the origin. For any non-negative measurable function $f$ on $\mathbb{R}^{+}$, we write
@@
In [1998BROWNIAN, Section 4], Aldous derived the law of the reflected Brownian bridge on time interval $[0,1]$ conditioned on having local time profile $\ell$, where $\ell$ may lie in a broad class of functions. This class includes, in particular, all continuous profiles. In our notation, this is exactly the law of $W^{\ell}$ for continuous profiles $\ell$ satisfying $h(\ell)<\infty$ and $I(\ell)=1$.
We now explain how this conditional law is described. Let $\mathcal{T}^{\ell}$ denote the rooted ordered real tree encoded by $W^{\ell}$, and let $\mathcal{T}^{\ell}_{k}$ denote its uniform $k$-point subtree. Aldous [1998BROWNIAN] and, subsequently, Warren [warren1999result] determined the law of $\mathcal{T}^{\ell}_{k}$, which will be described in detail later in Section 2.3.2. Since the family $\{\mathcal{T}^{\ell}_{k}\}_{k\geq 1}$ is consistent in $k$, the correspondence theorem of Aldous [1993Aldous, Theorem 15] [^6] A direct extension of the correspondence theorem from excursion functions to bridge functions applies here. yields a version of $W^{\ell}$. Moreover, this version of the conditional law is continuous in $\ell$ in a suitable sense.
Next, we extend Aldous’ version of $W^{\ell}$ to all continuous profiles $\ell$ with $h(\ell)<\infty$ (i.e., without the restriction $I(\ell)=1$). This follows from the scaling invariance of reflected Brownian motion and its local times:
@@
$\tilde{\ell}(s):=I(\ell)^{-1/2}\,\ell\big(I(\ell)s\big),\qquad s\geq 0,$
@@
Many properties for the case $I(\ell)=1$ (cf. [1998BROWNIAN, Construction 1]) extend to this general setting, including continuity in $\ell$ and the following:
@@
    For any continuous profile $\ell$ with $h(\ell)<\infty$, the local time profile of $W^{\ell}$ is $\ell$; that is, for every non-negative measurable $f$ on $\mathbb{R}_{+}$,
    $\int_{0}^{\infty}\ell(s)\,f(s)\,\mathrm{d}s\ =\ \int_{0}^{I(\ell)}f\big(W^{\ell}_{t}\big)\,\mathrm{d}t\quad\text{a.s.}.$ (8)
@@
In [warren98], Warren and Yor provided an explicit construction of $W^{\ell}$ for continuous profiles $\ell$ satisfying $\ell(0)>0$ and $h(\ell)<\infty$. Roughly speaking, they introduced a process called the Brownian burglar, which is obtained through a suitable time-space transformation of a Brownian path. A version of $W^{\ell}$ can then be constructed from the Brownian burglar by applying another time-space transformation.
In particular, this version of $W^{\ell}$ is continuous in $\ell$ in the following sense (cf. [li2024inverting, Proposition 5.5]): if $(\ell^{n})$ and $\ell$ are continuous profiles with $h(\ell_{n})\rightarrow h(\ell)$ and $\ell_{n}\rightarrow\ell$ uniformly on compact subsets of $(0,h(\ell))$, then
$W^{\ell^{n}}\xrightarrow{d}W^{\ell}\quad\text{in }C[0,\infty),$
@@
By comparing the continuity of the canonical version $W^{\ell}$ with respect to $\ell$, one can readily deduce that the two versions coincide for all continuous profiles $\ell$ with $h(\ell)<\infty$. Hence, the Warren–Yor’s approach gives an explicit realization of $W^{\ell}$.

### v2

@@ 1.4 Organization of the paper
The remainder of the paper is organized as follows. In Section 2, we present several important preliminaries, including the formal definition of Cannings trees, Aldous’ theorem on the convergence of contour functions, and the construction of the conditioned Brownian motion $W^{\ell}$. Section 3 introduces the coalescent process and derives some of its basic properties, including the CDFI property. Section 4 is devoted to proving the tightness of the rescaled trees. Finally, in Section 5, we complete the proof of the main theorem by establishing the finite-dimensional convergence and the joint convergence of the contour and height functions. Appendix A provides a counterexample demonstrating the near-sharpness of condition (H2a).
@@ 2.1 Cannings tree
where by convention $(\mathbb{N}_{+})^{0}:=\{\emptyset\}$. Every rooted ordered tree can be canonically represented as a prefix-stable finite subset of $\mathcal{U}$. See, for example, [8, Section 1.1] for details.
Using this formalism, we now give a precise definition of the Cannings tree associated with an inhomogeneous Cannings model. Every $u\in\mathcal{U}$ can be represented as $u=(u_{1},\ldots,u_{m})\in({\mathbb{N}}_{+})^{m}$ for some $m\in{\mathbb{N}}$, with the convention that $m=0$ corresponds to $u=\emptyset$; we write $|u|=m$ for its length. For $u\in\mathcal{U}$ and $j\in\mathbb{N}_{+}$, we denote by $uj$ the concatenated sequence $(u_{1},\ldots,u_{m},j)$. For $u,v\in\mathcal{U}$, we write $u\prec v$ if $u$ precedes $v$ in the lexicographical order on $\mathcal{U}$.
@@
Intuitively, this construction starts by introducing an artificial ancestor, the root node $\emptyset$, which serves as the common parent of all individuals in generation $1$. Each vertex at height $s$ (i.e., with label length $s$) represents an individual in generation $s$, and the corresponding entry $\nu^{s}_{i}$ in the offspring vector determines the number of children of the $i$-th individual (in lexicographic order) in generation $s$. These children are then represented by appending $1,2,\ldots,\nu^{s}_{i}$ to the parent’s representation. The tree structure constructed in this way encodes the entire genealogy of the population.
@@ 2.2 Foundations of tree encoding and convergence
This section introduces the framework for encoding trees by height and contour functions, and then provides necessary and sufficient conditions for the convergence of contour functions.
@@ 2.2.1 Geometric trees
In this section, we extend the notion of rooted ordered trees by assigning a positive length to each edge, and thereby introduce the concept of a rooted ordered geometric tree. This enriched structure is better suited for studying convergence, as it allows for the rescaling of edge lengths.
@@
The function $w$ naturally induces a metric on $V(T)$ defined by
$d(v,v^{\prime}):=\sum_{e\in[\![v,v^{\prime}]\!]}w(e),\quad v,v^{\prime}\in V(T),$
where $v,v^{\prime}]\!]$ denotes the unique path connecting $v$ and $v^{\prime}$ in $T$. In particular, with $\rho$ denoting the root of $T$,
@@
The height function $H(T)$ of a rooted ordered tree $(T,w)$ records the height of each vertex according to the lexicographic order on $T$. Specifically, we can rewrite
$V(T)=\{v_{0},v_{1},\dots,v_{k-1}\}\quad\text{with}\quad\rho=v_{0}\prec v_{1}\prec\cdots\prec v_{k-1}.$
The height function of $(T,w)$ is defined by
@@ 2.2.2 Encoding compact real trees by bridge functions
Following [8, Section 2.1], compact real trees are defined as compact metric spaces satisfying the standard tree properties. A canonical way to construct such trees, which simultaneously endows them with a root and an ordered structure, is through encoding functions drawn from the space
@@
Specifically, given $g\in C_{\rm bridge}$, we define a pseudo-metric
$d_{g}(s,t):=g(s)+g(t)-2\inf_{r\in[s\wedge t,s\vee t]}g(r),\quad\text{for }s,t\geq 0,$
which induces an equivalence relation:
@@
The quotient space $\mathcal{T}_{g}:=[0,1]/\sim$, equipped with the inherited metric $d_{g}$, is a compact real tree by [8, Theorem 2.2]. Let $p_{g}:[0,1]\to\mathcal{T}_{g}$ be the canonical projection. We regard $\mathcal{T}_{g}$ as rooted at $p_{g}(0)$ and equip it with an order by declaring $s_{1}\prec_{g}s_{2}$ if and only if
@@
Le Gall’s encoding of compact real trees in [8] uses continuous functions with compact support. Our encoding through $C_{\rm bridge}$ functions is essentially equivalent because, in Le Gall’s setting, the function $g(x)$ and its rescaled version $g(cx)$ encode the same real tree. Thus, without loss of generality, we can always assume $g\in C_{\rm bridge}$.
@@
where $x,y]\!]\subset\mathcal{T}_{g}$ denotes the unique geodesic path between $x,y\in\mathcal{T}_{g}$. This subtree is naturally viewed as a rooted ordered geometric tree: its vertex set consists of the root $\rho$, the sampled points $p_{g}(t_{1}),\dots,p_{g}(t_{k})$, and all branch points along their geodesics; edges are the geodesic segments connecting consecutive vertices, endowed with their metric lengths; and the order is inherited from $\preceq_{g}$.
@@
Suppose $g,\widetilde{g}\in C_{\rm bridge}$ differ only by a strictly increasing time change; that is, there exists a strictly increasing homeomorphism $\phi:[0,1]\to[0,1]$ such that $g(t)=\widetilde{g}\big(\phi(t)\big)$ for all $t\in[0,1].$ Then the map
$\Psi:\mathcal{T}_{g}\longrightarrow\mathcal{T}_{\widetilde{g}},\qquad\Psi\big(p_{g}(t)\big):=p_{\widetilde{g}}\big(\phi(t)\big),$
defines an isomorphism between the rooted ordered real trees $\mathcal{T}_{g}$ and $\mathcal{T}_{\widetilde{g}}$. In particular, for any $t_{1},\ldots,t_{k}\in[0,1]$, the subtrees $\mathcal{T}_{g}(t_{1},\ldots,t_{k})\quad\text{and}\quad\mathcal{T}_{\widetilde{g}}\big(\phi({t_{1}}),\ldots,\phi({t_{k}})\big)$ are isomorphic.
@@ 2.2.3 Convergence of contour functions
where $b^{i}_{(0)}=b^{i}_{(k)}$ is the root of $T^{i}$, and for $1\leq j\leq k-1$, $b^{i}_{(j)}$ denotes the branch point of $V^{i}_{(j)}$ and $V^{i}_{(j+1)}$ in $T^{i}$. It is straightforward to check that $d$ induces a metric on rooted ordered geometric trees with $k$ distinguished vertices, modulo rooted order-preserving isometries.
@@
Let $(\mathcal{T},d)=(\mathcal{T}_{g},d_{g})$ be a random rooted ordered real tree encoded by a random $g\in C_{\rm bridge}$. Consider a sequence $(\mathscr{T}^{n}:n\geq 1)$ of random ordered discrete trees with $a_{n}:=|V(\mathscr{T}^{n})|\to\infty$, and suppose that the maximal edge length of $\mathscr{T}^{n}$ converges to 0 in distribution. Then the following statements are equivalent:
@@
    For each $1\leq k\leq a_{n}$, let $(V^{n}_{1},\dots,V^{n}_{a_{n}})$ be a uniform random ordering of the vertices of $\mathscr{T}^{n}$, and denote by $\mathscr{T}^{n}_{k}$ the rooted ordered subtree spanned by the root $\rho$ and $\{V^{n}_{1},\dots,V^{n}_{k}\}$. Then it holds that:
    • Tightness: for any $\delta>0$,
      $\lim_{k\to\infty}\limsup_{n\to\infty}\mathbb{P}(\Delta(n,k)>\delta)=0,$
      where
      $\Delta(n,k):=\max_{v\in V(\mathscr{T}^{n})}\min_{u\in V(\mathscr{T}^{n}_{k})}d(v,u).$ (2.2)
    • Finite-dimensional convergence: for every fixed $k\in\mathbb{N}_{+}$,
      $\big(\mathscr{T}^{n}_{k},\{V^{n}_{1},\dots,V^{n}_{k}\}\big)\overset{d}{\longrightarrow}\big(\mathcal{T}_{g,k},\{p_{g}(U_{1}),\dots,p_{g}(U_{k})\}\big)$ (2.3)
      under the distance (2.1), where $\{V^{n}_{1},\dots,V^{n}_{k}\}$ and $\{p_{g}(U_{1}),\dots,p_{g}(U_{k})\}$ are considered as multisets.
@@
Aldous originally proved this equivalence in [2, Theorem 20] for the case where the limiting tree is a compact binary real tree encoded by an excursion. In that setting, his notion of finite-dimensional distributions differs from (2.3). Nevertheless, Theorem 2.6 is already implicit in Aldous’ proof. Indeed, the first step of his proof shows that his definition is equivalent to requiring that, for every $k\geq 1$,
$\begin{aligned} &\big(C_{U^{n}_{(1)}}(\mathscr{T}^{n}),\inf_{U^{n}_{(1)}<t<U^{n}_{(2)}}C_{t}(\mathscr{T}^{n}),C_{U^{n}_{(2)}}(\mathscr{T}^{n}),\inf_{U^{n}_{(2)}<t<U^{n}_{(3)}}C_{t}(\mathscr{T}^{n}),\ldots,C_{U^{n}_{(k)}}(\mathscr{T}^{n})\big)\\ &\overset{d}{\longrightarrow}\big(g(U_{(1)}),\inf_{U_{(1)}<t<U_{(2)}}g(t),g(U_{(2)}),\inf_{U_{(2)}<t<U_{(3)}}g(t),\ldots,g(U_{(k)})\big),\end{aligned}$ (2.4)
where $(U^{n}_{1},\ldots,U^{n}_{k})$ is a random vector sampled uniformly without replacement from $\{0,\ldots,2(a_{n}-1)\}$, with $U^{n}_{(i)}$ denoting its order statistics, and $U_{(1)},\ldots,U_{(k)}$ are the order statistics of $k$ i.i.d. uniform $(0,1)$ random variables. By [2, Lemma 12], (2.4) is readily seen to be equivalent to (2.3). This is the only step in his argument that relies on the assumption of a binary limiting tree; the remainder of the proof proceeds without invoking it. In our setting, as the root of the limiting tree is not necessarily binary, the formulation in (2.3) is better suited.
@@ 2.3 Conditioned Brownian motion and its subtree distribution
Recall the notion of a continuous profile introduced in Section 1. In this section, we first construct a canonical version of $W^{\ell}$ based on Aldous’ construction [3]. We then study the laws of the uniform $k$-point subtrees induced by $W^{\ell}$ and by the time-changed process $W^{\ell^{\sigma}}_{\alpha^{-1}(t)}$. These results will be used in the proof of finite-dimensional convergence in Section 5.1.
@@ 2.3.1 Canonical version of $W^{\ell}$
Recall that $L_{t}(x)$ denotes the local time of $W$ at level $x$ up to time $t$, and that $\tau_{u}$ represents the inverse local time at the origin. For any non-negative measurable function $f$ on $\mathbb{R}_{+}$, we write
@@
In [3, Section 4], Aldous derived the law of the reflected Brownian bridge on the time interval $[0,1]$ conditioned on its local time profile $L_{1}=\ell$, where $\ell$ may lie in a broad class of functions that includes all continuous profiles. In our notation, this is exactly the law of $W^{\ell}$ for continuous profiles $\ell$ satisfying $h(\ell)<\infty$ and $I(\ell)=1$.
We now explain how this conditional law is described. Let $\mathcal{T}^{\ell}$ denote the rooted ordered real tree encoded by $W^{\ell}$, and let $\mathcal{T}^{\ell}_{k}$ denote its uniform $k$-point subtree. Aldous [3] and, subsequently, Warren [18] determined the law of $\mathcal{T}^{\ell}_{k}$, which will be described in detail later in Section 2.3.2. Since the family $\{\mathcal{T}^{\ell}_{k}\}_{k\geq 1}$ is consistent in $k$, the correspondence theorem of Aldous [2, Theorem 15] [^4] A direct extension of the correspondence theorem from excursion functions to bridge functions applies here. yields a version of $W^{\ell}$. Moreover, this version of the conditional law is continuous in $\ell$ in a suitable sense as shown in [3, Construction 1].
Next, we extend Aldous’ version of $W^{\ell}$ to all continuous profiles $\ell$ with $h(\ell)<\infty$ (i.e., without the restriction $I(\ell)=1$). This follows from the scaling invariance of reflected Brownian motion and its local times.
@@
$\tilde{\ell}(s):=I(\ell)^{-\frac{1}{2}}\,\ell\big(I(\ell)^{\frac{1}{2}}s\big),\qquad s\geq 0,$
@@
Many properties that hold in the case $I(\ell)=1$ (cf. [3, Construction 1]) extend to this general setting, including continuity in $\ell$ and the following:
@@
    For any continuous profile $\ell$ with $h(\ell)<\infty$, the local time profile of $W^{\ell}$ is $\ell$; that is, a.s. for every non-negative measurable $f$ on $\mathbb{R}_{+}$,
    $\int_{0}^{I(\ell)}f\big(W^{\ell}_{t}\big)\,\mathrm{d}t\ =\ \int_{0}^{\infty}\ell(s)\,f(s)\,\mathrm{d}s.$ (2.5)
@@
In [20], Warren and Yor provided an explicit construction of $W^{\ell}$ for continuous profiles $\ell$ satisfying $\ell(0)>0$ and $h(\ell)<\infty$. Roughly speaking, they introduced a process called Brownian burglar, which is obtained through a suitable time-space transformation of a Brownian path. A version of $W^{\ell}$ can then be constructed from the Brownian burglar by applying another time-space transformation.
In particular, this version of $W^{\ell}$ is continuous in $\ell$ in the following sense (cf. [10, Proposition 5.5]): if $(\ell^{n})$ and $\ell$ are continuous profiles with $h(\ell_{n})\rightarrow h(\ell)$ and $\ell_{n}\rightarrow\ell$ uniformly on compact subsets of $(0,h(\ell))$, then
$W^{\ell^{n}}\overset{d}{\longrightarrow}\,W^{\ell}\qquad\text{in }C[0,\infty),$
@@
By comparing the continuity of the canonical version $W^{\ell}$ with respect to $\ell$, one can readily deduce that the two versions coincide for all continuous profiles $\ell$ with $h(\ell)<\infty$. Hence, the Warren–Yor’s construction gives an explicit realization of $W^{\ell}$.